# Cold Jupiters and small planets: friends, foes, or indifferent?

### v1

@@
Finding out whether there is any correlation between the presence of short-period small planets (SPs) with $P\lesssim 100$ d ($a\lesssim 0.4$ AU) and $1<M_{\rm p}<20\,\hbox{$\mathrm{M}_{\oplus}$}$ and that of outer cold Jupiters (CJs) with $a=1-10$ AU and $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ around solar-type stars may provide crucial constraints on the models of formation and/or migration of SPs. However, somehow discrepant results about the occurrence rates of CJs in SP systems have been reported in the literature, with a few recent works claiming a strong SP-CJ correlation but only at super-solar metallicity and/or mass of the host stars. Here we homogeneously recomputed the occurrence rates of CJs at average, sub-solar ($\rm[Fe/H]<-0.1$), solar ($\rm-0.1\leq[Fe/H]\leq 0.1$), and super-solar ($\rm[Fe/H]>0.1$) metallicity as well as at average and sub-intervals of stellar mass, namely 0.6-0.8, 0.8-1.0, and 1.0-1.2 $\mathrm{M}_{\odot}$, with (i) a carefully-selected sample of 217 SP systems detected through transits or radial velocity, and (ii) a large comparison sample of 1167 solar-type stars, regardless of the possible presence of SPs. We determined the integrated occurrence rate of CJs in SP systems to be $f_{\rm CJ|SP}=11.1^{+2.5}_{-1.8}\%$ at average metallicity $\overline{\rm[Fe/H]}=-0.011\pm 0.005$ and mass $\overline{M_{\star}}=0.916\pm 0.012\,\hbox{$\mathrm{M}_{\odot}$}$; this is consistent with the estimated frequencies of CJs in both the comparison sample ($f_{\rm CJ}=9.8^{+0.9}_{-0.8}\%$ at $\overline{\rm[Fe/H]}=-0.072\pm 0.009$ and $\overline{M_{\star}}=0.994\pm 0.004\,\hbox{$\mathrm{M}_{\odot}$}$) and the HARPS-N survey of transiting SP systems. We found a possible correlation ($f_{\rm CJ|SP}>f_{\rm CJ}$) only at super-solar mass and metallicity, namely $1.0\leq M_{\star}<1.2\,\hbox{$\mathrm{M}_{\odot}$}$ and $\rm[Fe/H]>0.1$, though with a statistical confidence less than $3\sigma$. To test some theoretical predictions, we also searched for possible SP-CJ relations as a function of SP and CJ multiplicity as well as SP composition, albeit with the inevitably limited current sample, and found none. We show that the architectures of SP systems are not indifferent to the presence of CJs as the multiplicity of SPs strongly depends on the CJ eccentricity, as expected from planetary dynamics. A more comprehensive understanding of the relation between SPs and CJs requires larger samples of SP systems. The increasing number of well-characterized systems and the anticipated discoveries from both the Gaia and PLATO missions will enable a definitive assessment of the impact of CJs on the formation of SPs.
@@ 1 Introduction
Whether the presence of short-period small planets (SPs) with $P\lesssim 100$ d (corresponding to semi-major axis of $a\lesssim 0.4$ AU) and $1<M_{\rm p}<20\,\hbox{$\mathrm{M}_{\oplus}$}$ around solar-type stars is somehow correlated with that of cold Jupiters (CJs), that is gaseous giant planets with $a\simeq 1-10$ AU and mass $0.5\leq M_{\rm p}\leq 20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ (or $0.3\leq M_{\rm p}\leq 13\,\hbox{$\mathrm{M}_{\rm Jup}$}$, depending on the definition), is becoming a very intriguing question in exoplanetary science because it could potentially reveal crucial features of planet formation and evolution. However, there are strong discrepancies about this possible correlation in the literature from both theoretical and observational perspectives.
As for the theory, an anti-correlation between SPs and CJs would be expected if CJs could either (i) act as a dynamical barrier to the inward migration of icy SPs from beyond the water snowline at $\sim 1-3$ AU (Izidoro et al. 2015) and/or (ii) reduce considerably the inward flux of pebbles to the inner regions of the protoplanetary disk, which is required to form dry SPs within the water snowline (Lambrechts et al. 2019). From the Generation 3 Bern simulations of planet formation and evolution, Schlecker et al. (2021) instead found no or a weak SP-CJ correlation, and predicted a strong architecture-composition link, with more icy (rocky) SPs in the absence (presence) of one or more CJs, due to the same above-mentioned effect of dynamical barrier by the CJ (Izidoro et al. 2015). Furthermore, Schlecker et al. (2021) predicted a scarcity of SPs at high stellar metallicity due to the disruption of inner planetary systems by more dynamically active CJs (see their Fig. 19). A recent work by Bitsch & Izidoro (2023) shows that it would also be possible to obtain a correlation between the presence of SPs and that of CJs, depending on the gas contraction rate for the formation of CJs: for slow gas contraction rates, the cores that form in the proximity of the water snowline are too small to effectively accrete large envelopes and can migrate inward thus becoming SPs, while cores at greater distances can grow into cold gas giants.
@@
The works by Zhu (2024) and BL24 were partly motivated by the fact that those authors noticed that several stars in the Kepler and K2 sample by B23 are metal poor, which might have biased the B23 estimate of $f_{\rm CJ|SP}$ towards lower values for the well-known relation between giant planets and stellar metallicity (e.g., Fischer & Valenti 2005; Johnson et al. 2010; Sousa et al. 2011; Mortier et al. 2013). However, the average metallicity of the B23 sample ($\overline{\rm[Fe/H]}=-0.066\pm 0.014$) and that of the Anglo Australian Telescope (AAT) RV survey ($\overline{\rm[Fe/H]}=-0.045\pm 0.018$) (Wittenmyer et al. 2020), which was used as a comparison sample by B23, are consistent within $1\sigma$ and are both lower than the peak of the metallicity at about $-0.1$ dex in a volume-limited sample of the solar neighborhood (e.g., Sousa et al. 2011).
Building upon the work of BL24, in the present article we recompute (i) $f_{\rm CJ|SP}$ with a larger and more carefully selected sample of SP systems, and (ii) $f_{\rm CJ}$ using three different blind RV campaigns, namely the AAT, CLS, and HARPS (Mayor et al. 2011; Trifonov et al. 2020) surveys, and merging the stars of these three surveys into a unique large sample. We then compare $f_{\rm CJ|SP}$ and $f_{\rm CJ}$ to search for possible SP-CJ correlations as a function of stellar metallicity and/or mass as well as planet properties, and draw our conclusions about their possible existence.
@@ 2.1 Small planet transit and radial velocity sample
The sample of SP systems to compute $f_{\rm CJ|SP}$ was built through queries to the NASA Exoplanet Archive [^3] https://exoplanetarchive.ipac.caltech.edu/ . on 2 December 2024 following the same criteria as BL24: the systems must have (1) at least one confirmed SP with $1<M_{\rm p}<20\,\hbox{$\mathrm{M}_{\oplus}$}$ [^4] Unlike BL24, we did not consider the criterion on the radius to be $R_{\rm p}\leq 4\,\hbox{$\mathrm{R}_{\oplus}$}$ as this information is not available for the RV planets. for transiting planets and $1<M_{\rm p}\sin{i}<20\,\hbox{$\mathrm{M}_{\oplus}$}$ for planets discovered by RV surveys; (2) a solar-type (i.e., FGK dwarf or sub-giant) star as host, which thus excludes M dwarfs with $M_{\star}\leq 0.6\,\hbox{$\mathrm{M}_{\odot}$}$, and (3) at least one publicly available RV dataset with more than 20 measurements and baseline $\Delta T\geq 1$ yr. However, unlike BL24, we initially discarded “mixed systems” with both SP and short-period ($P<100$ d) planets with $M_{\rm p}\geq 20\,\hbox{$\mathrm{M}_{\oplus}$}$, such as the WASP-47 system which also hosts a hot Jupiter (WASP-47b) in between the two SPs WASP-47e and d (e.g., Vanderburg et al. 2017; Nascimbeni et al. 2023); indeed, it is unclear whether the theoretical predictions mentioned in Sect. 1 apply to such systems in which giant planets, especially the hot Jupiters, may have undergone a substantial migration process (e.g., Dawson & Johnson 2018). Nonetheless, we computed the occurrence rates by also taking the mixed systems into account, and noted no significant differences (see Sect. 3). We cleaned the BL24 sample from a couple of outliers like TOI-1853b, which has $M_{\rm p}=73\pm 3>20\,\hbox{$\mathrm{M}_{\oplus}$}$ (Naponiello et al. 2023), or Kepler-22b, whose orbital period is $P=290>100$ d (Borucki et al. 2012).
Our SP sample results in 217 systems, of which 134 were detected by transit surveys and 83 with RVs. It is therefore considerably larger than that of BL24 (184 mixed systems), and contains 148 systems in common (see Table LABEL:table_SPsample). The inclusion of the mixed systems leads to 242 systems, 151 and 91 of which were found with transits and RVs, respectively (Tables LABEL:table_SPsample and LABEL:table_Mixed_sample). As in BL24, for each system we selected the available RV dataset with the best sensitivity to the detection of CJs (see Sect. 3).
@@ 2.2 Radial velocity comparison samples
The mass, metallicity and uncertainty on the metallicity of the selected solar-type stars from the three surveys were all taken from the Hypatia catalog [^5] https://www.hypatiacatalog.com/ . (Hinkel et al. 2014) and are given in Table LABEL:table_RVsurveys.
We also merged the three aforementioned RV surveys to create a large comparison sample with 1167 stars, 213 of which are in common between two or even all three surveys (Table LABEL:table_RVsurveys). In this way, we could compute a more precise and accurate $f_{\rm CJ}$ (see Tables 19 and 23).
2.3 Comparison of the small planet and the RV merged samples
The distributions of the stars of both the SP and RV merged samples with stellar metallicity and mass are compared in Figs. 4 and 5, respectively.
The Kolmogorov-Smirnov (K-S) test between the two empirical cumulative distribution functions (eCDF) of metallicity gives a relatively low p-value of 0.021 as the SP sample has a lower fraction of metal-poor stars at $\rm[Fe/H]\lesssim-0.2$ than the comparison sample (see Fig. 4).
@@ 2.3 Comparison of the small planet and the RV merged samples
In Fig. 6 we show the stellar metallicity distributions and eCDFs of both the SP and RV merged samples at the three intervals of stellar mass considered in Sect. 3, namely $0.6-0.8$, $0.8-1.0$, and $1.0-1.2\,\hbox{$\mathrm{M}_{\odot}$}$. While the eCDFs of metallicity are compatible in the $0.6-0.8\,\hbox{$\mathrm{M}_{\odot}$}$ and $0.8-1.0\,\hbox{$\mathrm{M}_{\odot}$}$ intervals with p-values of 99% and 27%, respectively, they considerably differ at super-solar mass (p-value $<1\%$) because of the considerably larger number of metal-poor stars in the RV merged sample (right panels of Fig. 6).
@@ 3 Statistical analyses
We determined $f_{\rm CJ|SP}$ from the SP sample (Sect. 2.1), and $f_{\rm CJ}$ from both the individual RV surveys and the comparison sample (Sect. 2.2), first at the average metallicity of the different samples, and then by considering three intervals in stellar metallicity, i.e. sub-solar ($\rm[Fe/H]<-0.1$), solar ($\rm-0.1\leq[Fe/H]\leq 0.1$), and super-solar ($\rm[Fe/H]>0.1$) metallicity. Similarly to BL24, we also split the samples in two broad intervals, namely $\rm[Fe/H]\leq 0$ and $\rm[Fe/H]>0$, even though, in our view, given the typical uncertainties of $\sim 0.04-0.07$ dex on $\rm[Fe/H]$, a star with $0.10.0<\rm[Fe/H]<$ 0.1 ($-0.1<\rm[Fe/H]<0.0$) should not be classified as metal rich (poor), but rather as solar in metallicity.
Given the different distributions of stellar mass in the SP and comparison samples (Sect. 2.3), we further considered three different intervals in stellar mass, namely $0.6\leq M_{\star}<0.8\,\hbox{$\mathrm{M}_{\odot}$}$, $0.8\leq M_{\star}<1.0\,\hbox{$\mathrm{M}_{\odot}$}$, and $1.0\leq M_{\star}<1.2\,\hbox{$\mathrm{M}_{\odot}$}$ [^6] BL25 also considered similar intervals of stellar mass to investigate the SP-CJ relation as a function of stellar metallicity and mass.. This is of interest because the frequency of giant planets also depends on stellar mass, though to a lesser extent than metallicity (e.g., Johnson et al. 2010; Fulton et al. 2021).
For each sample and its relative sub-samples in metallicity we used the binomial statistics
@@
where $d$ is the number of systems with at least one confirmed CJ, and $N_{\star,\rm eff}$ is the “effective” number of stars given by the product of the number of stars of the survey, $N_{\star}$, and the average survey sensitivity (or completeness) to CJs, $C$, i.e. $N_{\star,\rm eff}=N_{\star}\cdot C$.
@@
The completeness of each individual star, $C_{\rm i}$, was computed in the same way as B23 and BL24 (see their Sect. 3), that is through experiments of injection and recovery of CJ RV signals in a logarithmic grid of 30x30 $\Delta M_{\rm p}$-$\Delta a$ cells covering the ranges $0.3-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ and $1-10$ AU, respectively. Specifically, for each cell, we simulated 200 RV signals of CJs at the epochs of the RV dataset by randomly varying (i) $M_{\rm p}$ and $a$ within the cell bounds, (ii) $T_{\rm c}$ within the range of $P$ corresponding to $a$ and the stellar mass $M_{\star}$ (Tables LABEL:table_SPsample, LABEL:table_Mixed_sample and LABEL:table_RVsurveys) from Kepler’s third law; (iii) $\cos{i}$ from 0 to 1, where $i$ is the orbital inclination; and iv) the argument of periastron $\omega$ from 0 to $2\pi$, while drawing the orbital eccentricity $e$ from a $\beta$ distribution (Kipping 2013). We then shifted every RV point at time $t$ following a Gaussian distribution with mean and standard deviations equal to the RV value and its uncertainty, respectively. $C_{\rm i}$ was estimated as the recovery rate of the simulated signals (Keplerian orbits or linear and quadratic trends if $\Delta T$ was considerably shorter than the simulated CJ orbital period) across the whole grid using the $\Delta BIC$ criterion (see B23 and BL24). The average completeness $C$ of each sample (or sub-sample) was then simply obtained by the mean of the completenesses of the stars in that sample: $C=(1/N_{\star})\sum_{i=1}^{N_{\star}}{C_{\rm i}}$. The average completenesses of all the samples described in Sect. 2 are displayed in Fig. 7. We note a very good agreement when comparing our average completenesses of the SP RV and transit samples (upper panels in Fig. 7) with those shown in Fig. A.1 of BL24 (for $a=1-10$ AU).
@@
We considered two ranges of CJ mass [^7] Note that only minimum masses, $M_{\rm p}\sin{i}$, are known for the majority of CJs as their orbital inclination $i$ cannot be determined from RVs alone., namely $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ and $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ (for comparison with B23), but we mainly rely on the former because the latter suffers from incompleteness issues between 0.3 and 0.5 $\mathrm{M}_{\rm Jup}$ (Fig. 7). For each CJ mass range and the aforementioned bins of stellar metallicity and/or mass we give the number of solar-type stars ($N_{\star}$), the number of stars with confirmed CJs ($d$), the average completenesses ($C$), and CJ occurrence rates from (i) the SP and the large comparison samples in Tables 19 and 20, (ii) the SP transit, RV, and transit+RV samples as well as the SP and mixed transit, RV, and transit+RV samples in Table 21, and (iii) the AAT, CLS, and HARPS surveys in Table 23. Tables LABEL:table_CJs_SPsample, LABEL:table_CJs_SPsample_mixed, and LABEL:table_CJs_RVsurveys list the known CJs and their properties in the SP, SP and mixed, and comparison samples. Fig. 11 shows on the left (right) the architectures of the 27 (23) systems with CJs with $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ ($0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$) in the SP sample, respectively.

### v2

@@
Determining whether there is any correlation between the presence of short-period small planets (SPs) with $P\lesssim 100$ d ($a\lesssim 0.4$ AU) and $1<M_{\rm p}<20\,\hbox{$\mathrm{M}_{\oplus}$}$ and that of outer cold Jupiters (CJs) with $a=1-10$ AU and $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ around solar-type stars may provide crucial constraints on models of formation and/or migration of SPs. However, discrepant results regarding the occurrence rates of CJs in SP systems have been reported in the literature, with some recent studies suggesting a strong SP-CJ correlation, but only at super-solar metallicities and/or masses of the host stars. Here, we homogeneously recomputed the occurrence rates of CJs at average, sub-solar ($\rm[Fe/H]<-0.1$), solar ($\rm-0.1\leq[Fe/H]\leq 0.1$), and super-solar ($\rm[Fe/H]>0.1$) metallicities, as well as at average and subintervals of stellar mass, namely 0.6-0.8, 0.8-1.0, and 1.0-1.2 $\mathrm{M}_{\odot}$, using: (i) a carefully selected sample of 217 SP systems detected through transits or radial velocity, and (ii) a large comparison sample of 1167 solar-type stars, regardless of the possible presence of SPs. We determine the integrated occurrence rate of CJs in SP systems to be $f_{\rm CJ|SP}=11.1^{+2.5}_{-1.8}\%$ at average metallicity $\overline{\rm[Fe/H]}=-0.011\pm 0.005$ and mass $\overline{M_{\star}}=0.916\pm 0.012\,\hbox{$\mathrm{M}_{\odot}$}$; this is consistent with the estimated frequencies of CJs in both the comparison sample ($f_{\rm CJ}=9.8^{+0.9}_{-0.8}\%$ at $\overline{\rm[Fe/H]}=-0.072\pm 0.009$ and $\overline{M_{\star}}=0.994\pm 0.004\,\hbox{$\mathrm{M}_{\odot}$}$) and the HARPS-N survey of transiting SP systems. We find a possible correlation ($f_{\rm CJ|SP}>f_{\rm CJ}$) only at super-solar mass and metallicity, namely $1.0\leq M_{\star}<1.2\,\hbox{$\mathrm{M}_{\odot}$}$ and $\rm[Fe/H]>0.1$, although with statistical confidence of less than $3\sigma$. To test some theoretical predictions, we also searched for possible SP-CJ relations as a function of SP and CJ multiplicity, as well as SP composition, albeit with the inevitably limited current sample, and we found none. We show that the architectures of SP systems are not indifferent to the presence of CJs, because the multiplicity of SPs strongly depends on the CJ eccentricity, as expected from planetary dynamics. A more comprehensive understanding of the relation between SPs and CJs requires larger samples of SP systems. The increasing number of well-characterized systems and the anticipated discoveries from both the Gaia and PLATO missions will enable a definitive assessment of the impact of CJs on the formation of SPs.
@@ 1 Introduction
Whether the presence of short-period small planets (SPs) with $P\lesssim 100$ d (corresponding to a semimajor axis of $a\lesssim 0.4$ AU) and $1<M_{\rm p}<20\,\hbox{$\mathrm{M}_{\oplus}$}$ around solar-type stars correlates with that of cold Jupiters (CJs) – gaseous giant planets with $a\simeq 1-10$ AU and mass $0.5\leq M_{\rm p}\leq 20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ (or $0.3\leq M_{\rm p}\leq 13\,\hbox{$\mathrm{M}_{\rm Jup}$}$, depending on the definition) – is an intriguing question in exoplanetary science because it could potentially reveal crucial features of planet formation and evolution. However, there are strong discrepancies about this possible correlation in the literature from both theoretical and observational perspectives.
Theoretically, an anticorrelation between SPs and CJs is expected if CJs either act as a dynamical barrier to the inward migration of icy SPs from beyond the water snowline at $\sim 1-3$ AU (Izidoro et al. 2015), and/or if they considerably reduce the inward flux of pebbles to the inner regions of the protoplanetary disk, which is required to form dry SPs within the water snowline (Lambrechts et al. 2019). From the Generation 3 Bern simulations of planet formation and evolution, Schlecker et al. (2021) instead find no or a weak SP-CJ correlation and predict a strong architecture-composition link, with more icy (rocky) SPs in the absence (presence) of one or more CJs, due to the same dynamical barrier effect of the CJ (Izidoro et al. 2015). Furthermore, Schlecker et al. (2021) predict a scarcity of SPs at high stellar metallicity due to the disruption of inner planetary systems by more dynamically active CJs (see their Fig. 19). A recent study by Bitsch & Izidoro (2023) shows that a correlation between the presence of SPs and that of CJs is possible, depending on the gas contraction rate for the formation of CJs: for slow gas contraction rates, the cores that form in the proximity of the water snowline are too small to effectively accrete large envelopes and can migrate inward, becoming SPs, while cores at greater distances can grow into cold gas giants.
@@
The studies by Zhu (2024), BL24, and BL25 were partly motivated by the observation that several stars in the Kepler and K2 sample from B23 are metal-poor, which might bias the B23 estimate of $f_{\rm CJ|SP}$ toward lower values for the well-known relation between giant planets and stellar metallicity (e.g., Fischer & Valenti 2005; Johnson et al. 2010; Sousa et al. 2011; Mortier et al. 2013). However, the average metallicity of the B23 sample ($\overline{\rm[Fe/H]}=-0.066\pm 0.014$) and that of the Anglo Australian Telescope (AAT) RV survey ($\overline{\rm[Fe/H]}=-0.045\pm 0.018$) (Wittenmyer et al. 2020), which B23 used as a comparison sample, are consistent within $1\sigma$ and are both lower than the peak metallicity of approximately $-0.1$ dex in a volume-limited sample of the solar neighborhood (e.g., Sousa et al. 2011).
Based on the work of BL24, in the present study we recompute: (i) $f_{\rm CJ|SP}$ with a larger and more carefully selected sample of SP systems and (ii) $f_{\rm CJ}$, using three different blind RV campaigns, namely the AAT, CLS, and HARPS (Mayor et al. 2011; Trifonov et al. 2020) surveys, and merging the stars of these three surveys into a single large sample. We then compare $f_{\rm CJ|SP}$ and $f_{\rm CJ}$ to search for possible SP-CJ correlations as a function of stellar metallicity, mass, or planet properties and draw our conclusions about their possible existence.
@@ 2.1 Small planet transit and radial velocity sample
The sample of SP systems to compute $f_{\rm CJ|SP}$ was built through queries to the NASA Exoplanet Archive [^3] https://exoplanetarchive.ipac.caltech.edu/ . on 2 December 2024 following the same criteria as BL24. The systems must have (1) at least one confirmed SP with $1<M_{\rm p}<20\,\hbox{$\mathrm{M}_{\oplus}$}$ [^4] Unlike BL24, we did not consider the criterion $R_{\rm p}\leq 4\,\hbox{$\mathrm{R}_{\oplus}$}$ as this information is not available for the RV planets. for transiting planets and $1<M_{\rm p}\sin{i}<20\,\hbox{$\mathrm{M}_{\oplus}$}$ for planets discovered by RV surveys; (2) a solar-type (i.e., FGK dwarf or sub-giant) host star, thereby excluding M dwarfs with $M_{\star}\leq 0.6\,\hbox{$\mathrm{M}_{\odot}$}$; and (3) at least one publicly available RV dataset with more than 20 measurements and baseline $\Delta T\geq 1$ yr. However, unlike BL24, we initially discarded “mixed systems” containing both SP and short-period ($P<100$ d) planets with $M_{\rm p}\geq 20\,\hbox{$\mathrm{M}_{\oplus}$}$, such as the WASP-47 system, which also hosts a hot Jupiter (WASP-47b) between the two SPs WASP-47e and d (e.g., Vanderburg et al. 2017; Nascimbeni et al. 2023). It is unclear whether the theoretical predictions mentioned in Sect. 1 apply to such systems in which giant planets, especially the hot Jupiters, may have undergone a substantial migration process (e.g., Dawson & Johnson 2018). Nonetheless, we also computed the occurrence rates by taking the mixed systems into account and noted no significant differences (see Sect. 3). We removed a few outliers from the BL24 sample, such as TOI-1853b, which has $M_{\rm p}=73\pm 3>20\,\hbox{$\mathrm{M}_{\oplus}$}$ (Naponiello et al. 2023), and Kepler-22b, whose orbital period is $P=290>100$ d (Borucki et al. 2012).
Our SP sample comprises 217 systems, of which 134 were detected by transit surveys and 83 by RVs. It is therefore considerably larger than that of BL24 (184 mixed systems) and contains 148 systems in common (see Table 1 in Additional Tables). Including the mixed systems leads to 242 systems, 151 and 91 of which were found with transits and RVs, respectively (Tables 1 and 2 in Additional Tables). As in BL24, for each system we selected the available RV dataset with the best sensitivity for the detection of CJs (see Sect. 3).
@@ 2.2 Radial velocity comparison samples
The mass, metallicity, and uncertainty on the metallicity of the selected solar-type stars from the three surveys were all taken from the Hypatia catalog [^5] https://www.hypatiacatalog.com/ . (Hinkel et al. 2014) and are listed in Table 3 of Additional Tables.
We also merged the three aforementioned RV surveys to create a large comparison sample with 1167 stars, 213 of which are common to two or all three surveys. In this way, we could compute a more precise and accurate $f_{\rm CJ}$ (see Tables 13 and 17).
2.3 Comparison of the small planet and radial velocity merged samples
The distributions of the stellar metallicity and mass for both the SP and RV merged samples are compared in Fig. 4.
The Kolmogorov-Smirnov (K-S) test between the two empirical cumulative distribution functions (eCDF) of metallicity yields a relatively low p-value of 0.021, as the SP sample contains a lower fraction of metal-poor stars at $\rm[Fe/H]\lesssim-0.2$ than the comparison sample (see Fig. 4, left panel).
@@ 2.3 Comparison of the small planet and radial velocity merged samples
In Fig. 5 we show the stellar metallicity distributions and eCDFs of both the SP and RV merged samples for the three intervals of stellar mass considered in Sect. 3, namely $0.6-0.8$, $0.8-1.0$, and $1.0-1.2\,\hbox{$\mathrm{M}_{\odot}$}$. While the eCDFs of metallicity are compatible in the $0.6-0.8\,\hbox{$\mathrm{M}_{\odot}$}$ and $0.8-1.0\,\hbox{$\mathrm{M}_{\odot}$}$ intervals, with p-values of 99% and 27%, respectively, they differ considerably at super-solar mass (p-value $<1\%$) because of the considerably larger number of metal-poor stars in the RV merged sample (right panels of Fig. 5).
@@ 3 Statistical analyses
We determined $f_{\rm CJ|SP}$ from the SP sample (Sect. 2.1) and $f_{\rm CJ}$ from both the individual RV surveys and the comparison sample (Sect. 2.2), first at the average metallicity of the different samples, and then by considering three intervals in stellar metallicity, i.e., sub-solar ($\rm[Fe/H]<-0.1$), solar ($\rm-0.1\leq[Fe/H]\leq 0.1$), and super-solar ($\rm[Fe/H]>0.1$) metallicity. Similarly to BL24, we also split the samples into two broad intervals, namely $\rm[Fe/H]\leq 0$ and $\rm[Fe/H]>0$. However, given the typical uncertainties of $\sim 0.04-0.07$ dex on $\rm[Fe/H]$, a star with $0.10.0<\rm[Fe/H]<$ 0.1 ($-0.1<\rm[Fe/H]<0.0$) should be better classified as solar in metallicity rather than as metal-rich (poor).
Given the different distributions of stellar mass in the SP and comparison samples (Sect. 2.3), we further considered three different intervals in stellar mass, namely $0.6\leq M_{\star}<0.8\,\hbox{$\mathrm{M}_{\odot}$}$, $0.8\leq M_{\star}<1.0\,\hbox{$\mathrm{M}_{\odot}$}$, and $1.0\leq M_{\star}<1.2\,\hbox{$\mathrm{M}_{\odot}$}$ [^6] BL25 also considered similar intervals of stellar mass to investigate the SP-CJ relation as a function of stellar metallicity and mass.. This is of interest because the frequency of giant planets also depends on stellar mass, though to a lesser extent than on metallicity (e.g., Johnson et al. 2010; Fulton et al. 2021).
For each sample and its relative subsamples in metallicity, we used binomial statistics:
@@
where $d$ is the number of systems with at least one confirmed CJ, and $N_{\star,\rm eff}$ is the “effective” number of stars, given by the product of the number of stars in the survey, $N_{\star}$, and the average survey sensitivity (or completeness) to CJs, $C$, i.e., $N_{\star,\rm eff}=N_{\star}\cdot C$.
@@
The completeness of each individual star, $C_{\rm i}$, was computed in the same way as in B23 and BL24 (see their Sect. 3), that is, through injection and recovery experiments of CJ RV signals in a logarithmic grid of 30x30 $\Delta M_{\rm p}$-$\Delta a$ cells covering the ranges $0.3-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ and $1-10$ AU, respectively. Specifically, for each cell, we simulated 200 RV signals of CJs at the epochs of the RV dataset by randomly varying: (i) $M_{\rm p}$ and $a$ within the cell bounds, (ii) $T_{\rm c}$ within the range of $P$ corresponding to $a$ and the stellar mass $M_{\star}$ (Tables 1-3 in Additional Tables) from Kepler’s third law; (iii) $\cos{i}$ from 0 to 1, where $i$ is the orbital inclination; and iv) the argument of periastron $\omega$ from 0 to $2\pi$, while drawing the orbital eccentricity $e$ from a $\beta$ distribution (Kipping 2013). We then shifted every RV point at time $t$ following a Gaussian distribution with mean and standard deviation equal to the RV value and its uncertainty, respectively. We estimated $C_{\rm i}$ as the recovery rate of the simulated signals (Keplerian orbits or linear and quadratic trends if $\Delta T$ was considerably shorter than the simulated CJ orbital period) across the whole grid using the $\Delta BIC$ criterion (see B23 and BL24). The average completeness $C$ of each sample (or subsample) was then obtained by the mean of the completenesses of the stars in that sample: $C=(1/N_{\star})\sum_{i=1}^{N_{\star}}{C_{\rm i}}$. The average completenesses of all the samples described in Sect. 2 are displayed in Fig. 6. We note a very good agreement when comparing our average completenesses of the SP RV and transit samples (upper panels in Fig. 6) with those shown in Fig. A.1 of BL24 (for $a=1-10$ AU).
@@
We considered two ranges of CJ mass [^7] Note that only minimum masses, $M_{\rm p}\sin{i}$, are known for the majority of CJs as their orbital inclination $i$ cannot be determined from RVs alone., namely $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ and $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ (for comparison with B23), but we mainly rely on the former because the latter suffers from incompleteness issues between 0.3 and 0.5 $\mathrm{M}_{\rm Jup}$ (Fig. 6). For each CJ mass range and the aforementioned bins of stellar metallicity and/or mass, we provide the number of solar-type stars ($N_{\star}$), the number of stars with confirmed CJs ($d$), the average completenesses ($C$), and CJ occurrence rates from (i) the SP and the large comparison samples in Tables 13 and 14, (ii) the SP transit, RV, and transit+RV samples as well as the SP and mixed transit, RV, and transit+RV samples in Table 15, and (iii) the AAT, CLS, and HARPS surveys in Table 17. Tables 4, 5, and 6 in Additional Tables list the known CJs and their properties in the SP, mixed, and comparison samples. Figure 7 shows the architectures of systems in the SP sample: on the left, the 27 systems with CJs of mass $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$, and on the right, the 23 systems with CJs of mass $0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$.